Two meteoroids are 2.4×105 from Earth's center and moving at 2.1 . One is headed straight for Earth, while the other is on a path that will come within 8500 of Earth's center.Find the speed of the second meteoroid at its closest approach to Earth.
1
Expert's answer
2012-06-19T09:14:54-0400
https://upload.wikimedia.org/wikipedia/ru/math/8/0/f/80fa38486b470275c02b665928ec3cc3.png

G=
6,67259·10−11
M=2,4 × 105
R=8500

v=sqrt(6,67259·10−11 ×2,4 × 105/8500)=sqrt(0.2* 10−11)
v=1.41E-6